OpenStudy (fanduekisses):

If f(x)= 3x-2/2x+1 then f'(x) =

OpenStudy (fanduekisses):

\[f(x)=\frac{ 3x-2 }{ 2x+1 }\]

OpenStudy (fanduekisses):

Can I use the quotient rule?

OpenStudy (fanduekisses):

Oh wait no because the terms in the numerator and in the denominator aren;t being multiplied

geerky42 (geerky42):

Yes, use quotient rule.

OpenStudy (anonymous):

oh no quotient rule for this one for sure

OpenStudy (saifoo.khan):

SATELLIGHT!!!!

OpenStudy (fanduekisses):

oh ok let me try

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\]with

OpenStudy (anonymous):

whoa SAIFOOOOOOOOOO!

OpenStudy (andrewthecookie):

^^

OpenStudy (anonymous):

\[\huge \color{red}{\text{CELERITY IN THE HOUSE!}}\]

OpenStudy (andrewthecookie):

Mods these days cx

OpenStudy (anonymous):

actually i means celebrity but whatever

OpenStudy (fanduekisses):

Ok I get 7/(2x+1)^2 lol I can't believe I forgot how to do this for a moment ^_^ thanks

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\]with \[f(x)=3x-2,f'(x)=3,f(x)=2x+1,g'(x)=2\]

OpenStudy (saifoo.khan):

Celeb = the one with the smartscore of 0.1K. xD

OpenStudy (anonymous):

yeah you are right \[\frac{7}{(2x+1)^2}\] does it